E connect Managerial Accounting: FA18 BUAC 20B- ACCOUNTING HW Ch 9 Question 9 fof 9) 2.00 points DataSpan, Inc, automated its plant at the start of the current year and installed a flexible manufacturing system. The company is also evaluating its suppliers and moving toward Lean Production. Many adjustsment problems have been encountered, including problems relating to performance measurement. After much study, the company has decided data relating to these measures for the first four months of operations. to use the performance measures below, and it has gathered Month Throughput time (days) Delivery cycle time (days) Manufacturing cycle efficiency (MCE) Percentage of on-time deliveries Total sales (units) 80% 81% 86% 93% 10,550 10.420 10,400 10.500 Management has asked for your help in computing throughput time, delivery cycle time, and MCE. The following average times have been logged over the last four months Average per Month (in 0.5 0.5 0.6 0.3 0.5 0.4 0.3 0.7 Move time per unit Process time per unit Wait time per order before start of production Queue time per unit Inspection time per unit 9.6 8.0 50 4.0 4.1 3.6 2.8 2.0 04 08 0.7 0.8 Required 1-a. Compute the throughput time for each month. (Round your answers to 1 decimal place.) Throughput T
media%2F923%2F923e8afa-578b-4765-9b91-06
media%2Fcc2%2Fcc2970e6-3ddc-4eb5-8ed9-92
Answer #1

Part 1-a

Throughput time

Month 1

5.5

days

Month 2

5.3

days

Month 3

4.4

days

Month 4

3.8

days

Explanation:

1

2

3

4

Throughput time in days:

Process time

0.5

0.4

0.3

0.7

Inspection time

0.4

0.8

0.7

0.8

Move time

0.5

0.5

0.6

0.3

Queue time

4.1

3.6

2.8

2.0

Total throughput time

5.5

5.3

4.4

3.8

Part 1-b

Manufacturing cycle efficiency (MCE)

Process time ÷ Throughput time

Month 1

(0.5/5.5) = 9.1

%

Month 2

(0.4/5.3) = 7.5

%

Month 3

(0.3/4.4) = 6.8

%

Month 4

(0.7/3.8) = 18.4

%

Part 1-c

Throughput time

Month 1

15.1

days

Month 2

13.3

days

Month 3

9.4

days

Month 4

8.8

days

Explanation:

1

2

3

4

Delivery cycle time in days:

Wait time

9.6

8.0

5.0

4.0

Total throughput time

5.5

5.3

4.4

3.8

Total delivery cycle time

15.1

13.3

9.4

8.8

Part 3-a

Month 5

Throughput Time

1.8

days

Manufacturing cycle efficiency (MCE)

38.9

%

Part 3-b

Month 6

Throughput Time

1.0

days

Manufacturing cycle efficiency (MCE)

70.0

%

Explanation for 3-a and 3-b:

5

6

Throughput time in days

Process time

0.7

0.7

Inspection time

0.8

0

Move time

0.3

0.3

Queue time

0.0

0.0

Total throughput time

1.8

1.0

5

6

Manufacturing cycle efficiency (MCE):Process time ÷ Throughput time

38.9%

(0.7/1.8)

70.0%

(0.7/1.0)

Part 1-a

Throughput time

Month 1

5.5

days

Month 2

5.3

days

Month 3

4.4

days

Month 4

3.8

days

Explanation:

1

2

3

4

Throughput time in days:

Process time

0.5

0.4

0.3

0.7

Inspection time

0.4

0.8

0.7

0.8

Move time

0.5

0.5

0.6

0.3

Queue time

4.1

3.6

2.8

2.0

Total throughput time

5.5

5.3

4.4

3.8

Part 1-b

Manufacturing cycle efficiency (MCE)

Process time ÷ Throughput time

Month 1

(0.5/5.5) = 9.1

%

Month 2

(0.4/5.3) = 7.5

%

Month 3

(0.3/4.4) = 6.8

%

Month 4

(0.7/3.8) = 18.4

%

Part 1-c

Throughput time

Month 1

15.1

days

Month 2

13.3

days

Month 3

9.4

days

Month 4

8.8

days

Explanation:

1

2

3

4

Delivery cycle time in days:

Wait time

9.6

8.0

5.0

4.0

Total throughput time

5.5

5.3

4.4

3.8

Total delivery cycle time

15.1

13.3

9.4

8.8

Part 3-a

Month 5

Throughput Time

1.8

days

Manufacturing cycle efficiency (MCE)

38.9

%

Part 3-b

Month 6

Throughput Time

1.0

days

Manufacturing cycle efficiency (MCE)

70.0

%

Explanation for 3-a and 3-b:

5

6

Throughput time in days

Process time

0.7

0.7

Inspection time

0.8

0

Move time

0.3

0.3

Queue time

0.0

0.0

Total throughput time

1.8

1.0

5

6

Manufacturing cycle efficiency (MCE):Process time ÷ Throughput time

38.9%

(0.7/1.8)

70.0%

(0.7/1.0)